# AFLNet Five Years Later: On Coverage-Guided Protocol Fuzzing

### v1

@@ 1 Introduction
However, finding vulnerabilities in protocol implementations is also difficult. First, a server is stateful and message-driven. It takes a sequence of messages (a.k.a. requests) from a client, processes the messages, and sends appropriate responses. Yet, the implemented protocol may not entirely correspond to the specified protocol, making model-based fuzzing approaches [1, 2] less effective. For instance, as shown in Figure 2, the Live555 streaming media server implements a state machine for the Real-Time Streaming Protocol (RTSP) that supports an unspecified transition between the INIT and PLAY states (shown in red). Second, the server’s response depends on both the current message and its internal state, which is influenced by earlier messages, posing challenges for vanilla coverage-guided greybox fuzzers like American Fuzzy Lop (AFL) and its extensions [3, 4].
@@
In 2020, motivated by the aforementioned challenges of stateful network protocol fuzzing and the pressing need for an effective tool by researchers and practitioners, we introduced AFLnet [6]—the first code- and state-coverage-guided greybox fuzzer. However, an extended technical discussion and a full empirical evaluation of all its components in a full-length article have since been outstanding.
@@ 2.1 Motivating Example: File Transfer Protocol (FTP)
We begin with an informal introduction of the main concepts behind server communication and the terminology we are using in this paper. A server is a software system that can be accessed remotely, e.g., via the Internet. A client is a software system that uses the services which are provided by a server. In order for the client to use the services of a server, the server must first establish a connection via a communication channel. A network socket is an endpoint for sending or receiving data and can be identified by an IP address and a port.
In order to exchange information, both network participants send messages. A message is a distinct data packet. A message sequence is a vector of messages. A valid order of messages is governed by a protocol. The protocol provides strict rules and regulations that determine how data is transmitted and ensures reliable communication between client and server. A message from the client is also called request while a message from the server is called response. Each request may advance the server state, e.g., from initial state to authenticated. The server state is a specific status of the server in the communication with the client.
@@ 2.2 Difficulties of Traditional Fuzzing Approaches
To find vulnerabilities in stateful protocols, there are several challenges for state-of-the-art fuzzing approaches, like coverage-based greybox fuzzing (CGF) [15, 16] and stateful blackbox fuzzing (SBF) [1, 17]. CGF is an effective automated vulnerability detection technique. It leverages code coverage information, obtained by lightweight code instrumentation, to retain and prioritize interesting seeds (e.g., input files) generated by mutation operators (e.g., bit flips and splicing) in an evolutionary fashion. However, vanilla CGF, like AFL and its extensions [18, 4, 3], neither know the server state information nor the required structure or order of the messages to be sent. These CGF were mainly designed to test stateless programs (e.g., file processing programs) which always produce the same output for the same input. No state is maintained or taken into account.
@@ 3.1 Recording and Replay for Fuzzing
In order to facilitate mutational fuzzing for message sequences, we first need to develop the capability to record and replay message sequences. In order to record a realistic message exchange between client and server, a network sniffer can be used. A network sniffer captures network traffic for a specified period of time. For instance, we can use tcpdump [^3] https://www.tcpdump.org/pcap.html to capture the traffic from a user-generated FTP session. The sniffer records the entire network traffic that can be filtered automatically. The relevant message exchange can be extracted using a packet analyzer. A packet analyzer can identify and distinguish different message exchanges between different nodes in the network. For instance, we used the packet analyzer Wireshark [^4] https://www.wireshark.org/ to automatically extract the sequence of FTP requests.
@@ 3.2 Lightweight Protocol Learning
The directed labelled graph which reconciles all state transitions that have been observed throughout the fuzzing campaign, we call as implemented protocol state machine (IPSM). Each node represents a state. Each directed edge represents a state transition. The edge is annotated with a request and response message. If there is an edge between two states $s_{1}$ and $s_{2}$, and the server is currently in state $s_{1}$ and receives a request that matches the one in the edge label, then the server sends a response that matches the one in the edge label and transitions to state $s_{2}$. An example illustrating the utility of the IPSM is shown in Figure 2.
@@ 3.4 Mutating a Message Sequence
51      AFLnet is mutation-based fuzzer, i.e., a seed message sequence is chosen from a corpus and mutated to generate new sequences. There are several advantages over existing generation-based approaches which generate new message sequences from scratch. First, a mutation-based approach can leverage a valid trace of real network traffic to generate new sequences that are likely valid—albeit entirely without a protocol specification. In contrast, a generation-based approach [19, 2, 17] requires a detailed protocol specification, including concrete message templates and the protocol state machine. Hence, BooFuzz [17], a generation-based approach, does not discover the unspecified state transition in Figure 2 in page 2. Second, a mutation-based approach allows to evolve a corpus of particularly interesting message sequences. Generated sequences that have led to the discovery of new states, state transitions, or program branches are added to the corpus for further fuzzing. This evolutionary approach is the secret sauce of the tremendous recent success of coverage-based greybox fuzzing.
@@ 3.5 Selecting Interesting Message Sequences
59      To achieve this, AFLnet shifts the code coverage to the left of the bitmap by a certain number of elements SHIFT_SIZE (i.e., STATE_SIZE $\times$ STATE_SIZE, where STATE_SIZE is the maximum number of states observed) and computes the map index of code branches using Equation 2.
60
@@
61      Following this adjustment, AFLnet reserves the bitmap space with SHIFT_SIZE many elements to state coverage. While observing a state transition $S_{1}\rightarrow S_{2}$, AFLnet computes the map index for this state transition using Equation 3, where prev_state and cur_state are state keys by numbering raw states $S_{1}$ and $S_{2}$ starting from the number 1, and then increments the value at the corresponding index by 1.
62
$\begin{gathered}map\_index=(prev\_state\times{\rm STATE\_SIZE}+cur\_state)\\ ~{}\%~{}{\rm MAP\_SIZE}\end{gathered}$ (3)
@@ 4 Implementation
66
Figure 4: Architecture and Implementation of Stateful Greybox Fuzzing into AFLnet.
@@
70      The Target State Selector takes information from the IPSM to select that state which AFLnet should focus on next. AFL implements the seed corpus (here, containing message sequences) as a linked list of queue entries. A queue entry is the data structure containing pertinent information about the seed input. In addition, AFLnet maintains a state corpus which consists of (i) a list of state entries, i.e., a data structure containing pertinent state information, and (ii) a hashmap which maps a state identifier to a list of queue entries exercising the state corresponding to the state identifier. Both, the target state selector and the Sequence Selector leverage the state corpus. The Sequence Mutator augments AFL’s fuzz_one method with protocol-aware mutation operators.
@@
73      Now we illustrate how all these components of AFLnet work together to fuzz the LightFTP server. Suppose AFLnet starts with only one pcap file containing the network traffic as shown in 1. First, Request Sequence Parser parses the pcap file to generate a single sequence (as visualized in Figure 3) and save it into the corpus $C$. At the same time, State Machine Learning constructs the initial IPSM based on the response codes; this initial IPSM contains black nodes and transitions in Figure 5. Suppose that Target State Selector selects state 331 (“USER foo OK”) as the target state, Sequence Selector will then randomly select a sequence from the sequence corpus $C$, which contains only one sequence at this moment. Afterward, Sequence Mutators identifies the sequence prefix (“USER foo” request), the candidate subsequence (“PASS foo” request), and the remaining subsequence as the suffix. By mutating the candidate subsequence using stacked mutators, Sequence Mutators may generate a wrong password request (“PASS bar”) leading to an error state (530 Not logged in). Following this wrong password, it replays the suffix (e.g., “MKD demo”, “CWD demo”) leading to a loop in the state 530 because all these commands are not allowed before successful authentication. Finally, the “QUIT” request is sent, and the server exits. Since the generated test sequence (as visualized in Figure 6) covers new state and state transitions (as highlighted in red in Figure 5), it is added into the corpus $C$ and the IPSM.
@@ 5 Evaluation
80      Experimental Infrastructure. All experiments were conducted on an Intel® Xeon® Platinum 8468V CPU with 192 logical cores clocked at 2.70GHz, 512GB of memory, and running Ubuntu 22.04.3 LTS. Each experiment runs for 24 hours. We report the average results over 10 runs to mitigate the impact of randomness.
@@
TABLE I: Average branch coverage and average state coverage across 10 runs of 24 hours achieved by AFLnetQueue compared to AFLnetCode.
@@
AFLnetQueue AFLnetCode Improv ${\boldsymbol{\hat{A}_{12}}}$ AFLnetQueue AFLnetCode Improv ${\boldsymbol{\hat{A}_{12}}}$
Bftpd 491.5 484.8 +1.38% 0.72 334.0 170.5 +0.96$\times$ 1.00
DCMTK 3086.3 3076.6 +0.32% 0.56 3.0 3.0 0.00$\times$ 0.50
DNSmasq 1217.6 1221.0 -0.28% 0.43 27364.0 282.5 +95.85$\times$ 1.00
Exim 2862.7 2847.7 +0.53% 0.46 75.7 61.6 +0.23$\times$ 1.00
forked-daapd 2401.3 2384.9 +0.69% 0.54 37.7 22.0 +0.71$\times$ 1.00
Kamailio 9752.4 9800.2 -0.49% 0.43 300.3 89.2 +2.37$\times$ 1.00
lightftp 347.2 355.8 -2.42% 0.21 388.7 179.0 +1.17$\times$ 1.00
Live555 2818.5 2809.8 +0.31% 0.53 87.9 75.1 +0.17$\times$ 1.00
OpenSSH 3300.0 3336.7 -1.10% 0.20 30480.9 93.5 +325.00$\times$ 1.00
ProFTPD 5309.6 5296.4 +0.25% 0.55 473.5 250.6 +0.89$\times$ 1.00
Pure-FTPd 1277.1 1268.0 +0.72% 0.45 420.2 292.1 +0.44$\times$ 1.00
TinyDTLS 575.7 574.4 +0.23% 0.50 37.5 30.5 +0.23$\times$ 1.00
@@
Both AFLnetQueue and AFLnetCode select interesting seeds based on the test order in the corpus queue. We did not include the original AFLnet for comparison, because it uses an interleaving strategy between the seed-queue order and the state heuristics to select interesting seeds. Since this experiment specifically focuses on the impact of additional state guidance, we developed the invariant fuzzer AFLnetQueue to eliminate the influence of seed-selection strategies.
92      Table I shows the average number of code branches and states covered by AFLnetQueue and AFLnetCode across all subjects. To quantify the improvement of AFLnetQueue over AFLnetCode, we report the percentage improvements in terms of code coverage and state coverage achieved in 24 hours, respectively (Improv), as well as the possibility that a random campaign of AFLnetQueue outperforms a random campaign of AFLnetCode (${{\hat{A}_{12}}}$). We consider the Vargha-Delaney effect size ${{\hat{A}_{12}}}\geq 0.71$ or ${{\hat{A}_{12}}}\leq 0.29$ to indicate a substantial advantage of AFLnetQueue over AFLnetCode, or vice versa.
@@
TABLE II: Average branch coverage across 10 runs of 24 hours achieved by AFLnet compared to AFLnetQueue and AFLnetIPSM.
@@
TABLE III: Average state coverage across 10 runs of 24 hours achieved by AFLnet compared to AFLnetQueue and AFLnetIPSM.
@@
97      These experimental results demonstrate that additional state feedback can effectively guide the fuzzer to explore more states. Unfortunately, the additional state guidance did show significant effectiveness in improving the code coverage, while it also has no obvious harm on most subjects (i.e., 10 out of 12 subjects). A possible explanation is that AFLnet considers response codes from a server as the representation of states; however, this may not be a good state definition for each subject, as noted by the follow-up works of AFLnet [28, 26].
@@
103      Table II shows the average code branches covered by AFLnet, AFLnetQueue and AFLnetIPSM across 10 runs of 24 hours. In addition, we report the improvement in the code coverage of AFLnet compared to AFLnetQueue and AFLnetIPSM, respectively (Improv), and the Vargha-Delaney effect size (${\hat{A}_{12}}$). Compared to the seed-selection strategy based on queue order (i.e.,AFLnetQueue), the interleaving seed-selection strategy (i.e., AFLnet) performed significantly better in some subjects (i.e., DCMTK, Kamailio, OpenSSH, and TinyDTLS). However, it underperformed in the subjects DNSmasq, forked-daapd, and Pure-FTPd. In the remaining subjects, both seed-selection strategies had similar performance. Compared to the strategy based on the state heuristics (i.e., AFLnetIPSM), the interleaving strategy consistently performed better although the improvement was not statistically significant in some subjects (e.g., Bftpd).
104      Table III shows the state coverage of AFLnet, AFLnetQueue and AFLnetIPSM in a similar format to our previous table. Overall, AFLnet outperformed both baseline tools AFLnetQueue and AFLnetIPSM in terms of state coverage. AFLnet covered 5.77% more states than AFLnetQueue, and 12.77% more states than AFLnetIPSM on average. Although AFLnet covers fewer states than AFLnetIPSM or AFLnetQueue in the subjects Bftpd, lightftp, and TinyDTLS, this underperformance is not significant given the values of ${{\hat{A}_{12}}}$.

### v3

@@ 1 Introduction
However, finding vulnerabilities in protocol implementations is also difficult. First, a server is stateful and message-driven. It takes a sequence of messages (a.k.a. requests) from a client, processes the messages, and sends appropriate responses. Yet, the implemented protocol may not entirely correspond to the specified protocol, making model-based fuzzing approaches [1, 2] less effective. For instance, as shown in Figure 2, the Live555 streaming media server implements a state machine for the Real-Time Streaming Protocol (RTSP) that unintentionally introduces an unspecified transition between the INIT and PLAY states (shown in red). Second, the server’s response depends on both the current message and its internal state, which is influenced by earlier messages, posing challenges for vanilla coverage-guided greybox fuzzers like American Fuzzy Lop (AFL) and its extensions [3, 4].
@@
In 2020, motivated by the aforementioned challenges of stateful network protocol fuzzing and the pressing need for an effective tool by researchers and practitioners, we introduced AFLnet [6]–the first code- and state-coverage-guided greybox fuzzer. However, an extended technical discussion and a full empirical evaluation of all its components in a full-length article have since been outstanding.
@@ 2.1 Motivating Example: File Transfer Protocol (FTP)
We begin with an informal introduction of the main concepts behind server communication and the terminology we are using in this paper. A server is a software system that can be accessed remotely, e.g., via the Internet. A client is a software system that uses the services which are provided by a server. In order for the client to use the services of a server, the client must first establish a connection via a communication channel. A network socket is an endpoint for sending or receiving data and can be identified by an IP address and a port.
In order to exchange information, both network participants send messages. A message is a distinct data packet. A message sequence is a vector of messages. A valid order of messages is governed by a protocol. The protocol provides strict rules and regulations that determine how data is transmitted and ensures reliable communication between client and server. A message from the client is also called request while a message from the server is called response [^3] In some protocols, e.g., mutual authentication in TLS, when the client is authenticating its identity to the server, the request comes from the server.. Each request may advance the server state, e.g., from initial state to authenticated. The server state is a specific status of the server in the communication with the client.
@@
TABLE I: Limitations of traditional fuzzing approaches in finding vulnerabilities in stateful protocols
Approach Description Representative Tools Limitations
Coverage-guided greybox fuzzing (CGF) Leverage code coverage information to retain and prioritize interesting seeds generated by mutation operators AFL [13] and its extensions  [14, 4, 3] Neither know the server state information nor the required message structure or order to be sent
Workaround solutions of coverage-guided greybox fuzzing (1) Write test harness for unit testing of specific program states of the server under test (SUT);
(2) Concatenate message structures into files and use them as seeds to do normal mutational file fuzzing For (1), libFuzzer [15];
For (2), AFL [13] For (1), cannot thoroughly test the interactions / transitions between several program states; requires substantial manual effort to write a correct test harness; and cannot test the whole server whose source code is not available;
For (2), inefficiency and ineffectiveness in bug finding due to no knowledge of which message to focus on and no state transition information
Stateful blackbox fuzzing (SBF) Traverse the given protocol model and leverage data models/grammars of messages to generate message sequences from scratch beSTORM [16], BooFuzz [17], Peach [1] and Sulley [18] Writing state models and data models involves much manual effort and expertise, which are also often error-prone; and learn nothing from past fuzzing execution
@@ 2.2 Difficulties of Traditional Fuzzing Approaches
To find vulnerabilities in stateful protocols, there are several challenges for state-of-the-art fuzzing approaches, like coverage-based greybox fuzzing (CGF) [13, 15] and stateful blackbox fuzzing (SBF) [1, 17]. The details of each approach and their corresponding limitations are listed in Section 2.1. CGF is an effective automated vulnerability detection technique. It leverages code coverage information, obtained by lightweight code instrumentation, to retain and prioritize interesting seeds (e.g., input files) generated by mutation operators (e.g., bit flips and splicing) in an evolutionary fashion. However, vanilla CGF, like AFL and its extensions [14, 4, 3], neither know the server state information nor the required structure or order of the messages to be sent. These CGF were mainly designed to test stateless programs (e.g., file processing programs) which always produce the same output for the same input. No state is maintained or taken into account.
@@ 3.1 Recording and Replay for Fuzzing
In order to facilitate mutational fuzzing for message sequences, we first need to develop the capability to record and replay message sequences. In order to record a realistic message exchange between the client and the server, a network sniffer can be used. A network sniffer captures network traffic for a specified period of time. For instance, we can use tcpdump [^4] https://www.tcpdump.org/pcap.html to capture the traffic from a user-generated FTP session. The sniffer records the entire network traffic that can be filtered automatically. The relevant message exchange can be extracted using a packet analyzer. A packet analyzer can identify and distinguish different message exchanges between different nodes in the network. For instance, we used the packet analyzer Wireshark [^5] https://www.wireshark.org/ to automatically extract the sequence of FTP requests.
@@ 3.2 Lightweight Protocol Learning
We refer to the directed labelled graph which reconciles all state transitions that have been observed throughout the fuzzing campaign as the implemented protocol state machine (IPSM). Each node represents a state. Each directed edge represents a state transition. The edge is annotated with a request and response message. If there is an edge between two states $s_{1}$ and $s_{2}$, and the server is currently in state $s_{1}$ and receives a request that matches the one in the edge label, then the server sends a response that matches the one in the edge label and transitions to state $s_{2}$. An example illustrating the utility of the IPSM is shown in Figure 2.
@@ 3.4 Mutating a Message Sequence
51      AFLnet is mutation-based fuzzer, i.e., a seed message sequence is chosen from a corpus and mutated to generate new sequences. There are several advantages over existing generation-based approaches which generate new message sequences from scratch. First, a mutation-based approach can leverage a valid trace of real network traffic to generate new sequences that are likely valid—albeit entirely without a protocol specification. In contrast, a generation-based approach [18, 2, 17] requires a detailed protocol specification, including concrete message templates and the protocol state machine. Hence, BooFuzz [17], a generation-based approach, does not discover the unspecified state transition in Figure 2 in page 2. Second, a mutation-based approach allows the fuzzer to evolve a corpus of particularly interesting message sequences. Generated sequences that have led to the discovery of new states, state transitions, or program branches are added to the corpus for further fuzzing. This evolutionary approach is the secret sauce of the tremendous recent success of coverage-based greybox fuzzing.
@@ 3.5 Selecting Interesting Message Sequences
59
Figure 4: Architecture and Implementation of Stateful Greybox Fuzzing in AFLnet.
60      To achieve this, AFLnet shifts the code coverage to the left of the bitmap by a certain number of elements SHIFT_SIZE, where SHIFT_SIZE is the number of the bitmap entries reserved to store state coverage information). The map index for code branches is then computed using Equation 2.
61
@@
62      Following this adjustment, AFLnet reserves the bitmap space with SHIFT_SIZE many elements to state coverage. While observing a state transition $S_{1}\rightarrow S_{2}$, AFLnet computes the map index for this state transition using Equation 3, where prev_state and cur_state are state keys by numbering raw states $S_{1}$ and $S_{2}$ starting from the number 1, and STATE_SIZE is the maximum number of states expected to be observed at the end of fuzzing campaign. The value of the corresponding index is then incremented by 1.
63
$\begin{gathered}map\_index=(prev\_state\times{\rm STATE\_SIZE}+cur\_state)\\ ~{}\%~{}{\rm SHIFT\_SIZE}\end{gathered}$ (3)
@@ 4 Implementation
70      The Target State Selector takes information from the IPSM to select the state that AFLnet should focus on next. AFL implements the seed corpus (here, containing message sequences) as a linked list of queue entries. A queue entry is the data structure containing pertinent information about the seed input. In addition, AFLnet maintains a state corpus which consists of (i) a list of state entries, i.e., a data structure containing pertinent state information, and (ii) a hashmap which maps a state identifier to a list of queue entries exercising the state corresponding to the state identifier. Both the target state selector and the Sequence Selector leverage the state corpus. The Sequence Mutator augments AFL’s fuzz_one method with protocol-aware mutation operators.
@@
73      Now we illustrate how all these components of AFLnet work together to fuzz the LightFTP server. Suppose AFLnet starts with only one pcap file containing the network traffic as shown in 1. First, Request Sequence Parser parses the pcap file to generate a single sequence (as visualized in Figure 3) and saves it into the corpus $C$. At the same time, State Machine Learning constructs the initial IPSM based on the response codes; this initial IPSM contains black nodes and transitions in Figure 5. Suppose that Target State Selector selects state 331 (“USER foo OK”) as the target state. Sequence Selector will then randomly select a sequence from the sequence corpus $C$, which contains only one sequence at this moment. Afterward, Sequence Mutators identifies the sequence prefix (“USER foo” request), the candidate subsequence (“PASS foo” request), and the remaining subsequence as the suffix. By mutating the candidate subsequence using stacked mutators, Sequence Mutators may generate a wrong password request (“PASS bar”) leading to an error state (530 Not logged in). Following this wrong password, it replays the suffix (e.g., “MKD demo”, “CWD demo”) leading to a loop in the state 530 because all these commands are not allowed before successful authentication. Finally, the “QUIT” request is sent, and the server exits. Since the generated test sequence (as visualized in Figure 6) covers new state and state transitions (as highlighted in red in Figure 5), it is added into the corpus $C$ and the IPSM.
@@ 5 Evaluation
80      Experimental Configuration and Infrastructure. We conducted all the experiments using the Git commit 62d63a5 of AFLnet. The SHIFT_SIZE parameter was set to half of the bitmap size, reserving one-half of the bitmap space for state coverage and the other half for code coverage. We set STATE_SIZE to 256 based on our observations from our preliminary experiments. All experiments were run on an Intel® Xeon® Platinum 8468V CPU with 192 logical cores clocked at 2.70GHz, 512GB of memory, and running Ubuntu 22.04.3 LTS. Each experiment runs for 24 hours. We report the average results over 10 runs to mitigate the impact of randomness.
@@
TABLE II: Average branch coverage and average state coverage across 10 runs of 24 hours achieved by AFLnetQueue compared to AFLnetCode.
@@
AFLnetCode AFLnetQueue Improv ${\boldsymbol{\hat{A}_{12}}}$ AFLnetCode AFLnetQueue Improv ${\boldsymbol{\hat{A}_{12}}}$
Bftpd 484.8 491.5 +1.38% 0.72 170.5 334.0 +0.96$\times$ 1.00
DCMTK 3076.6 3086.3 +0.32% 0.56 3.0 3.0 0.00$\times$ 0.50
DNSmasq 1221.0 1217.6 -0.28% 0.43 282.5 27364.0 +95.85$\times$ 1.00
Exim 2847.7 2862.7 +0.53% 0.46 61.6 75.7 +0.23$\times$ 1.00
forked-daapd 2384.9 2401.3 +0.69% 0.54 22.0 37.7 +0.71$\times$ 1.00
Kamailio 9800.2 9752.4 -0.49% 0.43 89.2 300.3 +2.37$\times$ 1.00
LightFTP 355.8 347.2 -2.42% 0.21 179.0 388.7 +1.17$\times$ 1.00
Live555 2809.8 2818.5 +0.31% 0.53 75.1 87.9 +0.17$\times$ 1.00
OpenSSH 3336.7 3300.0 -1.10% 0.20 93.5 30480.9 +325.00$\times$ 1.00
ProFTPD 5296.4 5309.6 +0.25% 0.55 250.6 473.5 +0.89$\times$ 1.00
Pure-FTPd 1268.0 1277.1 +0.72% 0.45 292.1 420.2 +0.44$\times$ 1.00
TinyDTLS 574.4 575.7 +0.23% 0.50 30.5 37.5 +0.23$\times$ 1.00
@@
Both AFLnetQueue and AFLnetCode select interesting seeds based on the test order in the corpus queue. We did not include the original AFLnet for comparison because it uses an interleaving strategy between the seed-queue order and the state heuristics to select interesting seeds. Since this experiment specifically focuses on the impact of additional state guidance, we developed the invariant fuzzer AFLnetQueue to eliminate the influence of seed-selection strategies.
92      Table II shows the average number of code branches and states covered by AFLnetQueue and AFLnetCode across all subjects. To quantify the improvement of AFLnetQueue over AFLnetCode, we report the percentage improvements in terms of code coverage and state coverage achieved in 24 hours, respectively (Improv), as well as the possibility that a random campaign of AFLnetQueue outperforms a random campaign of AFLnetCode (${{\hat{A}_{12}}}$). We consider the Vargha-Delaney effect size ${{\hat{A}_{12}}}\geq 0.71$ or ${{\hat{A}_{12}}}\leq 0.29$ to indicate a substantial advantage of AFLnetQueue over AFLnetCode, or vice versa.
@@
TABLE III: Average branch coverage across 10 runs of 24 hours achieved by AFLnet compared to AFLnetQueue and AFLnetIPSM.
@@
TABLE IV: Average state coverage across 10 runs of 24 hours achieved by AFLnet compared to AFLnetQueue and AFLnetIPSM.
@@
97      These experimental results demonstrate that additional state feedback can effectively guide the fuzzer to explore more states. The additional state guidance did show significant effectiveness in improving the code coverage, and it also has no obvious harm on most subjects (i.e., 10 out of 12 subjects). A possible explanation is that AFLnet considers response codes from a server as the representation of states; however, this may not be a good state definition for each subject, as noted by the follow-up works of AFLnet [28, 26].
@@
103      Table III shows the average code branches covered by AFLnet, AFLnetQueue and AFLnetIPSM across 10 runs of 24 hours. In addition, we report the improvement in the code coverage of AFLnet compared to AFLnetQueue and AFLnetIPSM, respectively (Improv), and the Vargha-Delaney effect size (${\hat{A}_{12}}$). Compared to the seed-selection strategy based on queue order (i.e.,AFLnetQueue), the interleaving seed-selection strategy (i.e., AFLnet) performed significantly better in some subjects (i.e., DCMTK, Kamailio, OpenSSH, and TinyDTLS). However, it underperformed in the subjects DNSmasq, forked-daapd, and Pure-FTPd. In the remaining subjects, both seed-selection strategies had similar performance. Compared to the strategy based on the state heuristics (i.e., AFLnetIPSM), the interleaving strategy consistently performed better although the improvement was not statistically significant in some subjects (e.g., Bftpd).
104      Table IV shows the state coverage of AFLnet, AFLnetQueue and AFLnetIPSM in a similar format to our previous table. Overall, AFLnet outperformed both baseline tools AFLnetQueue and AFLnetIPSM in terms of state coverage. AFLnet covered 5.77% more states than AFLnetQueue, and 12.77% more states than AFLnetIPSM on average. Although AFLnet covers fewer states than AFLnetIPSM or AFLnetQueue in the subjects Bftpd, LightFTP, and TinyDTLS, this underperformance is not significant given the values of ${{\hat{A}_{12}}}$.